# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v2

@@ 5 Experimental Verifications
Experiment setup. We now verify our analysis in practical settings. We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023). We first fine-tune a GPT-2-large-774M model (Radford et al., 2019) on 55k samples of PKU-SafeRLHF-QA, and obtain the SFT model. We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle. All experiments are repeated for $3$ seeds. Please see Appendix C for more details.
Implementation details. For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset. For approximate optimization, we instead compute the empirical BT loss. We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$. During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$. Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
Verifications of Section 3. We train online DPO and RLHF on PKU-SafeRLHF-Prompt, following the practice of Dong et al. (2024); Shi et al. (2025). For the strong reward condition, we directly adopt the GPT2-large-harmless model as a perfectly-learned reward model. For the weak reward condition, we train the SFT model on PKU-SafeRLHF-safer by replacing the projection matrix with a linear head, freezing all layers except the linear head and the last block. For the strong policy condition, we fully train the SFT model, while for the weak policy condition, we freeze the first half of the blocks of the SFT model. Results are shown in Figures 2 and 3.
Verifications of Section 4. We train DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al. (2024). We train on two types of preference: “better” and “safer”, and down-sample the corresponding training datasets to 11k-99k samples. For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head. The models are trained under the same setting, and all achieve at least $85\%$ training accuracy. Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
@@ 6 Related Work
Here, we focus on comparing with the most relevant prior study, Nika et al. (2024). First, unlike Nika et al. (2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable upper bound of policy bias in the unregularized version. Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes. Third, we improve the statistical analysis of Nika et al. (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\tilde{\mathcal{O}}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
@@ Acknowledgement
SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P. Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship. RZ acknowledges the support of NSF TRIPODS II DMS-2023166. RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
@@ Appendix A Bonus: How Can We Better Model Reward From Preference Signals?
$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\phi}}^{\pi}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;\underset{y\sim\pi}{\mathbb{E}}[r_{\phi}(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;.$
If $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, this solution further admits the closed form $\pi_{\theta^{\star}(r_{\phi})}(y)=\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)/Z(\phi)$, where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)$ is the partition function. If the goal is to output a policy that performs well under the ground-truth reward $r^{\star}$, then reward learning should aim to find a model $r_{\phi}$ such that the resulting policy $\pi_{\theta^{\star}(r_{\phi})}$ maximizes the underlying “real” objective:
$r_{\phi^{\star}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmin}}\;\underbrace{-\beta\log Z(\phi)-\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:\mathcal{L}_{\textup{new}}(\phi)}\;.$
Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section B.11):
@@
which corresponds to the gradient of an $\ell_{2}$ distance of pairwise difference:
@@
Comparison with MLE. The reward model $r_{\phi}$ are typically learned via MLE from preference data, which does not consider the fact that the learned reward will ultimately be used to induce a policy. Now revisit the MLE objective:
@@
whose gradient is:
@@
Comparing Equation 4 with Equation 6, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 6 to further align with Equation 4. And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$. And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$. If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it can provide token-level reward information. However, it is worth noting that model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2. Therefore, when faced with a representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
@@ B.1 Proof of Proposition 1
Since $r^{\star}\in\mathcal{F}$, RLHF exactly recovers $r^{\star}$ during reward learning. The policy optimization stage then solves $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi}$, so by definition, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$.
@@ B.3 Proof of Proposition 3
Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$. DPO, instead, minimizes a proxy loss defined over pairwise preferences. Since $\pi_{\textup{DPO}}\in\Pi$, we have $V_{r^{\star}}^{\pi_{\textup{DPO}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}$, which proves the first claim.
@@ B.4 Proof of Proposition 5
Since $r^{\star}\notin\mathcal{F}$, RLHF recovers an approximation $\hat{r}\in\mathcal{F}$ via reward learning. It then computes a policy $\pi_{\textup{RLHF}}$ that maximizes $V_{\hat{r}}^{\pi}$ over $\Pi$. In general, this policy is sub-optimal under $r^{\star}$ (see Proposition 3), and thus $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}$.
DPO directly optimizes a preference-based loss over $\Pi$. Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V_{\Pi}^{\star}$.
@@
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}\;.$
@@ B.8 Numerical Proof of Proposition 4
By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$. Now we construct an environment under Condition 2, such that online DPO cannot outperform DPO, even with PILAF sampler. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward:

### v3

@@ 5 Experimental Verifications
5.1 Controlled experimental verifications
Experiment setup. We now verify our analysis in practical and controlled settings. We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023). We first fine-tune a GPT-2-large-774M model (Radford et al., 2019) on 55k samples of PKU-SafeRLHF-QA, and obtain the SFT model. We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle. All experiments are repeated for $3$ seeds. Please see Appendix C for more details.
Verifications of Section 3. We train online DPO and RLHF on PKU-SafeRLHF-Prompt, following the practice of Dong et al. (2024); Shi et al. (2025). For the strong reward condition, we directly adopt the GPT2-large-harmless model as a perfectly-learned reward model. For the weak reward condition, we train the SFT model on PKU-SafeRLHF-safer by replacing the projection matrix with a linear head, freezing all layers except the linear head and the last block. For the strong policy condition, we fully train the SFT model, while for the weak policy condition, we freeze the first half of the blocks of the SFT model. Results are shown in Figures 2 and 3. The empirical findings align closely with our theoretical predictions:
  • Figure 2 (Condition 1) aligns with Propositions 1 and 2: increasing the reward scale amplifies the second-order deviation in online DPO’s objective, causing larger deviation from the RLHF optimum, as our theory predicts.
  • Figure 3 (left, Condition 2) confirms Propositions 3 and 4: with a realizable reward model but restricted policy class, RLHF outperforms DPO.
  • Figure 3 (middle, Condition 3) confirms Proposition 5: with a mis-specified reward model but realizable policy class, DPO outperforms RLHF.
  • Figure 3 (right, Condition 4) exhibits behavior consistent with our double model mis-specification analysis: relative performance can depend on the comparative expressive power of $\mathcal{F}$ versus $\mathcal{F}_{\Pi}$. In our setup, the reward model is less expressive, leading RLHF to underperform.
Verifications of Section 4. We train DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al. (2024). We train on two types of preference: “better” and “safer”, and down-sample the corresponding training datasets to 11k-99k samples. For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head. The models are trained under the same setting, and all achieve at least $85\%$ training accuracy. Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks. This corroborates our theoretical separation result in Theorem 10: pure reward learning is statistically more sample-efficient than the surrogate reward learning performed by DPO.
@@ 5.1 Controlled experimental verifications
5.2 Connection with exisiting empirical studies.
Although our experiments are intentionally small-scale, our conclusions are supported by several recent large-scale empirical studies:
  • Table 1 of Razin et al. (2025b) show that DPO is worse than RLHF when learning the reward, supporting our claim in Section 4. Their models are Gemma-2-2B-IT, Qwen series, and Llama series, and the datasets are UltraFeedback and RewardMATH.
  • Figure 4 of Swamy et al. (2025) supports our observation in Section 3 that policy model requires larger scale to prevent mis-specification.
  • Figure 6 of Swamy et al. (2025) supports our Proposition 4 that online DPO cannot improve DPO much when the policy model is mis-specified. Their model is Pythia-1.4B.
  • Table 6 of Xu et al. (2024b) demonstrates that RLHF is stronger than DPO with a perfect reward signal, supporting our Proposition 3. Their models are Llama series, and benchmarks are SafeRLHF and APPS.
  • Table 5 of Ivison et al. (2024) demonstrates that RLHF can beat DPO under a good reward model, supporting our Proposition 3 and double-model-mis-specification analysis. Their models are Tulu2 series.
  • Figure 2 of Rafailov et al. (2023) demonstrates that DPO can beat PPO under a weak reward model, supporting our Proposition 5 and double-model-mis-specification analysis. Their reward model is initialized from GPT2-large model and trained for only $3$ epochs on the datasets, and is thus relatively weak.
These large-scale studies support the qualitative trends predicted in our theoretical analysis.
@@ 6 Related Work
Here, we focus on comparing with the most relevant prior study, Nika et al. (2024). First, unlike Nika et al. (2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable policy bias in the unregularized version. Second, we provide an analysis of model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n$ is large enough, and we are not limited to linear model classes. Third, we improve the statistical analysis of Nika et al. (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\tilde{\mathcal{O}}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
Concurrent works. We also acknowledge two concurrent works (Razin et al., 2026; Gopalan et al., 2026), which both show that DPO is not an idealized estimator for reward modeling. We differ in approaches: Razin et al. (2026) study learning dynamics, and Gopalan et al. (2026) propose a geometric characterization.
@@ Acknowledgement
SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P. Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship. RZ and MF acknowledge the support of NSF TRIPODS II DMS-2023166. The work of MF was supported in part by awards NSF CCF 2212261 and NSF CCF 2312775. RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
@@ Appendix A Bonus: How Can We Better Model Reward From Preference Signals?
In this section, we study the following relevant question:
@@
$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\phi}}^{\pi}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;\underset{y\sim\pi}{\mathbb{E}}[r_{\phi}(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;,$
and the reward model can further be projected to a surrogate reward model:
$\hat{r}_{\theta^{\star}(r_{\phi})}(y):=\beta\log\pi_{\theta^{\star}(r)}(y)/\pi_{{\mathsf{ref}}}(y)\;.$
As revealed in Appendix C of Ouyang et al. [2022] and Section 3.3 of Swamy et al. [2025], it is uncommon to deploy a reward model with a larger scale than the policy model. And thus to ensure practical relevance, we will focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$. Since $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, the solution further admits the closed form $\pi_{\theta^{\star}(r_{\phi})}(y)=\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)/Z(\phi)$, where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)$ is the partition function, and $\hat{r}_{\theta^{\star}(r_{\phi})}=r_{\phi}$ module constant shift.
If the goal is to output a policy that performs well under the ground-truth reward $r^{\star}$, then reward learning should aim to find a model $r_{\phi}$ such that the resulting policy $\pi_{\theta^{\star}(r_{\phi})}$ optimizing the underlying “real” objective:
$V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underbrace{\beta\log Z(\phi)+\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:-\mathcal{L}_{\textup{new}}(\phi)}\;.$
Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Appendix B.12):
@@
which corresponds to the gradient of an $\ell_{2}$ loss of pairwise difference:
@@
Therefore, to optimize the “real” reward model quality is to optimize an $\ell_{2}$ loss of pairwise difference:
$-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}{\mathbb{E}}\Delta^{2}_{y,y^{\prime}}(r^{\star}-r_{\phi})\;.$ (7)
Comparison with MLE. The reward model $r_{\phi}$ are typically learned via MLE from preference data, which does not consider the fact that the learned reward will ultimately be used to induce a policy. Let the distribution of the preference data be $\mu$ (by default $\mu$ is $\pi_{{\mathsf{ref}}}$, but can be any distribution here). Now revisit the MLE objective:
@@
whose gradient is (see detailed calculations in Appendix B.12):
@@
Following Equation 8, we can see that the gradient of DPO is
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\theta)\propto-\mathbb{E}_{y,y^{\prime}\sim\mu}[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))][\nabla(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))]\;,$
and the gradient of reward modeling is
$\nabla_{\phi}\mathcal{L}_{\text{RM}}(\phi)\propto-\mathbb{E}_{y,y^{\prime}\sim\mu}[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))][\nabla(r_{\phi}(y)-r_{\phi}(y^{\prime}))]\;.$
Comparing Equation 5 with Equation 8, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 8 to further align it with Equation 5. And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$. And this by-product explains why is PILAF sampler (a variant of online sampler, see Definition 1) introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$. If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it can provide token-level reward information. However, it is worth noting that model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2. Therefore, when faced with a representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
@@ B.1 Proof of Proposition 1
Since $r^{\star}\in\mathcal{F}$, RLHF exactly recovers $r^{\star}$ during reward learning. The policy optimization stage then solves $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi}$, so by definition, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}$.
@@ B.3 Proof of Proposition 3
Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}$. DPO, instead, minimizes a proxy loss defined over pairwise preferences. Since $\pi_{\textup{DPO}}\in\Pi$, we have $V_{r^{\star}}^{\pi_{\textup{DPO}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V^{\Pi}_{r^{\star}}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}$, which proves the first claim.
@@ B.4 Proof of Proposition 5
Since $r^{\star}\notin\mathcal{F}$, RLHF recovers an approximation $\hat{r}\in\mathcal{F}$ via reward learning. It then computes a policy $\pi_{\textup{RLHF}}$ that maximizes $V_{\hat{r}}^{\pi}$ over $\Pi$. In general, this policy is sub-optimal under $r^{\star}$ (see Proposition 3), and thus $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V^{\Pi}_{r^{\star}}$.
DPO directly optimizes a preference-based loss over $\Pi$. Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V^{\Pi}_{r^{\star}}$.
@@
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}\;.$
@@ B.8 Numerical Proof of Proposition 4
By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$. Now we construct an environment under Condition 2, such that online DPO cannot outperform DPO, even with PILAF sampler. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward: